

********************************************************************
UARS WEEKLY STATUS REPORT - 13 FEB 2004
********************************************************************
***********
SOLSTICE
***********
During this week SOLSTICE continued to operate in good health gathering
solar data on all available orbital tracking periods. On average, we
received 57% of each day's data amounting to about 439 minutes of solar
observing time per day.
***********
SUSIM
***********
SUSIM successfully gathered all of its daily mid (1.1 nm) resolution,
Mg II, occultation, solar pointing offset and high resolution
monitoring scans during this past week for periods of UARS real-time
telemetry or onboard tape recording.
***********
PEM
***********
Level 2 PEM HEPSA data files are being reprocessed for submission to the
DAAC. This data has been designated as version 2; however, there is no format
difference from the version 1 data format. Version 2 HEPSA data is being
generated because it was discovered that there exists processing errors within
some of the HEPSA version 1 data files. Thus, if you are using any HEPSA
version 1 data files from the DAAC, please replace them with the version 2
data files. Currently, HEPSA data between 1991274 and 2003149 has been
reprocessed and archived at the GSFC DAAC.
Level 2 PEM MEPS data file documentation is complete and submitted to
the DAAC. MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 1993209. MEPS data files between
1993210 and 1993216 are waiting for DAAC collection.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
This report covers 31 January 2004 (Orbit 67781, GMT Day 031) through
06 February 2004 (Orbit 67885, GMT Day 037).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 033/1910 67823 DAY-ONLY TO OFF
PEM 036/0017 67856 POWER ON - DAY ONLY
The beta angle ranged between the angles of 33.7 degrees to 40.8
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 21 February 2004 (DOY 052).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.96 to 1.28 degrees C. Additional
battery performance data for each battery on 30 January and 06
February 2004 are provided in Appendix D.
The clock error ranged from 12.1 to -11.0 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 10
SUSIM = 01
HALOE = 01
PEM = 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 01
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT completely backed up the operational CMS on 02
February 2004. The following day the CMS had the new 3-TDRS code
installed. Then on 04 February 2004 the FOT successfully executed the
uplink of the 3-TDRS patch. The uplink of the flight software went
extremely smoothly and the FOT would like to thank the FSW group for
all its hard work in making it so easy. The UARS Eclipse ground
system was also updated with the new build following the flight
software patch. This build contains new telemetry and commands to
support the 3-TDRS functionality. This transition went very well also
and all ground systems are in their 3-TDRS mode.
One problem was found during the transition however. The spacecraft
ephemeris was rejected onboard UARS late Wednesday evening. It was
found that the load was truncated and zero-filled for approximately
the last third of the load. The spacecraft continued to use the
existing ephemeris and was safe until a new load could be built and
uplinked. After some problems building this new load and a few tests
it was determined that the new CMS code was the source of the problem.
A new load was built from the old CMS code and uplinked. The FOT will
continue to build the spacecraft ephemeris on the old version of
software until the problem with the code can be resolved. The FOT
also has new checks in place to verify the ephemeris contents before
uplinking them to the spacecraft.
MEETINGS
The FOT called a working meeting on 05 February to review the CMS
problem and determine the future course of action.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 21 February 2004 (DOY 052).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 06 February 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
67781 031/0028 04-038 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
67799 032/0430 04-039 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
67801 032/0757 04-040 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
67812 033/0240 04-041 SUSIM CLEARED EXCEPTION MONITOR
67814 033/0532 04-042 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
67820 033/1408 04-043 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
67821 033/1628 04-044 HALOE STOW CMDS
67823 033/1910 04-045 PEM DAY-ONLY TO OFF
67836 034/1547 04-046 OBC DUMP TABLE 4 & 5
67840 034/2232 04-047 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
67841 035/0019 04-048 OBC ATC NOOPS 3 TDRSS
67853 035/1904 04-049 OBC 3 TDRSS OBC PATCH
67856 036/0017 04-050 PEM POWER ON - DAY ONLY
67858 036/0236 04-051 OBC DUMP OBC BANKS 0 THRU 15
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 06 February 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 06 February 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
67781 031/0034 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
67799 032/0432 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
67801 032/0806 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
67814 033/0534 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
67820 033/1408 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
67840 034/2232 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 06 February 2004
30 January 2004 - GMT Day 030 Beta = 33.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.31 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-179.2 +5.79/+4.50 +28.4/-14.4 25.1 1.04 20.9
3 +11.2/-123.2 +5.14/+3.55 +28.0/-15.2 25.1 1.03 20.7
06 February 2004 - GMT Day 037 Beta = 40.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.31 +00.0/+00.0 07.0 0.00 0.0
2 +16.8/-145.6 +5.47/+4.51 +26.8/-15.2 25.1 1.03 20.2
3 +11.2/-100.8 +4.51/+3.23 +26.4/-15.2 25.1 1.02 19.9
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period
APPENDIX E
UARS ANOMALY REPORTS
Week ending 06 February 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-009 FOT 67854 035/2126 N/A EPHEMERIS REJECTED
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-004 FOT 67589 018/0624 N/A IP NOCC switch problems
04-006 FOT 67684 024/1322 34874 HPA Failure
04-007 FOT 67688 024/2017 N/A Could Not Load
04-008 SOLSTICE 67688 024/2017 N/A SOLSTICE Detectors RED
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-001 FOT See List Jan 04 30832 Generic Late Acquisition
04-009 FOT 67854 035/2126 N/A EPHEMERIS REJECTED